OpenStudy (anonymous):

Brenda made two similar glass display boxes shaped like triangular prisms. The length of the sides of the base of the first display box is 4 feet, and the length of the sides of the base of the second display box is 8 feet. If the volume of the first display box is 42 ft3, what is the volume of the second display box?

OpenStudy (amistre64):

4^2 sin(60) ---------- = base triangle area i believe 2

OpenStudy (amistre64):

 base = sqrt(3) sqrt(3) * h = 42 h = 42/sqrt(3) , is that right?

OpenStudy (anonymous):

336 ft3 672 ft3 1344 ft3 2688 ft3

OpenStudy (anonymous):

is the choices

OpenStudy (amistre64):

then with any luck we will arrive at one of those choices :)

OpenStudy (anonymous):

ok lets do it

OpenStudy (amistre64):

 4*4 sin(60) ---------- = base area 2 4*4*sqrt(3) --------- = 4sqrt(3) for the base area 4 4sqrt(3) * h = 42 h = 21/2sqrt(3)

OpenStudy (amistre64):

if we double this to be similar to the other one in measure we get a height of 21/sqrt(3)

OpenStudy (amistre64):

base area of other one is: 8*8 sin(60) 8*8*sqrt(3) ---------- = ---------- = 16 sqrt(3) 2 4 16sqrt(3) * 21/sqrt(3) = 16*21 ... 210 +126 = 336 ... maybe?

OpenStudy (amistre64):

yay!! its an option, if its right or not tho ... lol

OpenStudy (anonymous):

good enough for me haha thanks

OpenStudy (amistre64):

yw, i cant see any fauxpauxes in my work so it might be good; at least an 83% chance of success lol